# Towards Domain Specification of Embedding Models in Medicine

### v1

@@ 5.2 Disease clustering
As previously demonstrated, our model outperforms all baselines on the clustering task (Table 4), confirming its unrivaled ability to group semantically related medical texts. Moreover, the qualitative t-SNE visualization in Figure 2—which projects Wikipedia-derived embeddings into two dimensions and color-codes each point by disease and body system—further corroborates this finding. Here, each disease category forms a distinct, tightly knit cluster with minimal overlap, illustrating the model’s remarkable sensitivity to fine-grained, disease-specific features. Together, these results not only reaffirm the model’s superior clustering capacity but also highlight its promise for downstream tasks such as disease stratification, clinical information retrieval, and diagnostic support.
Figure 2: t‑SNE Visualization of Wikipedia Disease Embeddings by Body System.
6 Code, Data and Model Availability
@@ 7 Conclusion
We introduced MedTE, a novel medical‐text embedding model trained exclusively via self‐supervised contrastive learning on a richly diverse corpus (PubMed abstracts, Wikipedia medical articles, ClinicalTrials.gov entries, MIMIC-IV notes, bioRxiv/medRxiv preprints, and large-scale medical QA). We also presented MedTEB, the first large-scale benchmark dedicated to medical embeddings, comprising 51 tasks spanning classification, clustering, pair classification, and retrieval. Across this suite, MedTE consistently outperforms both domain-specific baselines (e.g., BioBERT, ClinicalBERT, Med-BERT) and leading general-purpose models (e.g., E5, SBERT), achieving an average score of 0.578 versus 0.525 for the nearest competitor, and establishing a new state of the art in medical text representation.
@@ 7.1 Challenges
Despite these advances, the variability and complexity of medical language demand even broader, more diverse training data—especially for low-resource specialties and rare diseases—to capture nuanced semantic relationships. Moreover, domain shifts between sources (e.g., clinical notes versus scientific abstracts) can still degrade performance on niche tasks, highlighting the need for more finely granulated corpora and adaptive sampling strategies.
@@ 7.2 Future Work
Possible extensions could include: first, the development of a dedicated reranker trained on medical QA pairs to boost retrieval precision in downstream applications; and second, the integration of MedTE alongside this reranker within an end‑to‑end QA-RAG pipeline augmenting an Agentic AI, thereby enabling dynamic decisions about when and how to retrieve, rerank, and respond to complex medical queries in real time. Together, these directions could push the frontier of intelligent medical NLP toward more accurate, context‑aware clinical decision‑support systems.

### v2

@@ 5.2 Disease clustering
As previously demonstrated, our model outperforms all baselines on the clustering task (Table 4), confirming its unrivaled ability to group semantically related medical texts. Moreover, the qualitative t-SNE visualization in Figure 2, which projects Wikipedia-derived embeddings into two dimensions and color-codes each point by disease and body system, further corroborates this finding. Here, each disease category forms a distinct, tightly knit cluster with minimal overlap, illustrating the model’s remarkable sensitivity to fine-grained, disease-specific features. Together, these results not only reaffirm the model’s superior clustering capacity but also highlight its promise for downstream tasks such as disease stratification, clinical information retrieval, and diagnostic support.
Figure 2: t-SNE Visualization of Wikipedia Disease Embeddings by Body System.
6 Code, Data, and Model Availability
@@ 7 Conclusion
We introduced MedTE, a novel medical-text embedding model trained exclusively via self-supervised contrastive learning on a richly diverse corpus (PubMed abstracts, Wikipedia medical articles, ClinicalTrials.gov entries, MIMIC-IV notes, bioRxiv/medRxiv preprints, and large-scale medical QA). We also presented MedTEB, the first large-scale benchmark dedicated to medical embeddings, comprising 51 tasks spanning classification, clustering, pair classification, and retrieval. Across this suite, MedTE consistently outperforms both domain-specific baselines (e.g., BioBERT, ClinicalBERT, Med-BERT) and leading general-purpose models (e.g., E5, SBERT), achieving an average score of 0.578 versus 0.525 for the nearest competitor, and establishing a new state of the art in medical text representation.
@@ 7.1 Challenges
Despite these advances, the variability and complexity of medical language demand even broader, more diverse training data, especially for low-resource specialties and rare diseases, to capture nuanced semantic relationships. Moreover, domain shifts between sources (e.g., clinical notes versus scientific abstracts) can still degrade performance on niche tasks, highlighting the need for more finely granulated corpora and adaptive sampling strategies.
@@ 7.2 Future Work
Possible extensions could include: first, the development of a dedicated reranker trained on medical QA pairs to boost retrieval precision in downstream applications; and second, the integration of MedTE alongside this reranker within an end-to-end QA-RAG pipeline augmenting an Agentic AI, thereby enabling dynamic decisions about when and how to retrieve, rerank, and respond to complex medical queries in real time. Together, these directions could push the frontier of intelligent medical NLP toward more accurate, context-aware clinical decision-support systems.